# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v3

@@ 3.2.7. Comparison
In Figure 6, for sparsity, power consumption ranges from 30W to 400W, with inference speeds between 8.67 tokens/s and 92.5 tokens/s. The energy efficiency ranges from 0.069 tokens/J to 1.421 tokens/J. TF-MVP (yoo2023tf, ) (ASIC) achieves the lowest power consumption with 30.0167W and FlightLLM (zeng2024flightllm, ) (FPGA) achieves the highest throughput with batch size 1 (though it also applies quantization). And TF-MVP (ASIC) achieves the highest energy efficiency with 1.421 tokens/J.
@@
  • For GPUs, power consumption ranges from 70W to 400W, with inference speeds between 20.83 tokens/s and 50 tokens/s, situated in the middle right part of the figure. The energy efficiency ranges from 0.069 tokens/J to 0.434 token/J. Compared to CPUs, GPUs can achieve higher absolute inference speeds due to their high computing power and high bandwidth. However, the energy efficiency difference between CPUs is not significant.
  • For FPGAs, power consumption ranges from 45W to 155W, with inference speeds between 55 tokens/s and 92.5 tokens/s, also in the upper right part of the figure. The energy efficiency ranges from over 0.597 tokens/J to 1.32 tokens/J, which is much higher than GPUs and CPUs.
3.3. Fast Decoding
3.3.1. Overview
Traditional autoregressive decoding typically generates text token by token, choosing only the highest probability token at each step, known as greedy sampling. While this approach is simple and easy to implement, it may lead to a lack of diversity and creativity in the generated results. Another autoregressive method called nucleus sampling (or top-p sampling) (holtzman2019curious, ) considers multiple candidates during generation by setting a cumulative probability threshold p, allowing for sampling within a certain range. Although this method offers more diversity than greedy sampling, it still operates in a step-by-step generation manner. Currently, fast decoding techniques can be mainly divided into two categories: speculative decoding and skip layer.
@@ 3.3.1. Overview
Figure 7. Fast decoding.
Table 5. Fast decoding on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@
3.3.2. GPU
@@
Speculative Decoding. Speculative decoding (stern2018blockwise, ) is proposed to overcome the inherently sequential process in the autoregressive decoding of LLM. The essential decoding mechanism is to make predictions (i.e., draft tokens) parallelly for multiple time steps and then select the longest prefix verified by a scoring model as the final output. Lookahead decoding (lookahead, ) adopts the Guess-and-Verify paradigm as the whole decoding mechanism which generates the draft tokens by $n$-gram method and verifies the draft tokens during the forward at the same time. For Llama-7B and Llama-13B models with different datasets, it achieves 65.12-94.51 tokens/s and 56.01-90.05 tokens/s on a single NVIDIA A100-80GB GPU. To improve the acceptance rate while maintaining generation quality, many works focus on the scheme of generating draft tokens. Medusa (medusa, ) adds extra decoding heads to generate multiple subsequent tokens in parallel and uses a tree-based attention mechanism to construct multiple candidate continuations and verify them simultaneously in each decoding step. For Vicuna-7B and Vicuna-13B, it achieves 129.86 tokens/s and 98.54 tokens/s on a single NVIDIA A100-80GB GPU. EAGLE (eagle1, ; eagle2, ) selects a single transformer layer with the same configuration as LLM as the draft model to make predictions autoregressively and combines the feature and token embedding as the input of the draft model. During the verification phase on the target model, EAGLE chooses the tree-based attention mechanism similar to Lookahead and Medusa to ensure the correct relationship between the draft tokens. For Vicuna-7B, Vicuna-13B, Llama2-Chat-7B, and Llama-2-Chat-13B models, it achieves 139.95 tokens/s, 132.31 tokens/s, 133.98 tokens/s, and 156.80 tokens/s on a single NVIDIA A100-80GB GPU. Based on the basic paradigm of speculative decoding for prediction by draft models and verification by target models, some studies explore the optimizations on the modules in it. Ouroboros (zhao2024ouroboros, ) constructs a phrase candidate pool from the verification process of LLMs to provide candidates for the draft token generation of the draft model. Different from Medusa and EAGLE, Ouroboros uses the smaller LLMs (e.g., DeepSeek-7B) as the draft models for the target models (e.g., DeepSeek-34B). For Yi-34B, DeepSeek-34B and CodeLlama-34B models, it achieves 61.20 tokens/s, 41.00 tokens/s and 39.2 tokens/s on a single NVIDIA A100-80GB GPU. During the prediction phase, Sequoia (chen2024sequoia, ) introduces a dynamic programming algorithm and a hardware-aware tree optimizer to find the optimal tree structure based on the runtime features and the given hardware platform. During the verification phase, Sequoia uses a novel sampling and verification method that outperforms prior work across different decoding temperatures. For Llama2-7B and Llama2-13B models, it achieves 169.68 tokens/s and 149.20 tokens/s on a single NVIDIA A100-80GB GPU.
@@
The studies mentioned above all require the help of an auxiliary model (e.g., a single transformer layer in EAGLE, several Medusa heads in Medusa, or a smaller LLM like Llama2-7B in Ouroboros) or the statistical methods (e.g., $n$-gram in Lookahead) to generate the predicted tokens and then the target model is utilized to verify the predicted tokens. Some other studies (zhang2023draft, ; liu2024kangaroo, ; layerskip, ) also explore the LLM inference acceleration without the need of auxiliary models, called self-speculative decoding. Draft&Verify (zhang2023draft, ) generates draft tokens by selectively skipping certain intermediate layers of LLMs. Subsequently, the draft tokens will be verified in one forward pass. For the Llama2-13B model, it achieves 62.23 tokens/s on a single NVIDIA A100-40GB GPU. Kangaroo (liu2024kangaroo, ) adopts a fixed shallow sub-network of the LLM as the draft model, with the remaining layers serving as the target model. To enhance the representation ability of the draft model (i.e., the shallow sub-network), it trains an adapter module to follow the sub-network. For Vicuna-7B and Vicuna-13B models, it achieves 138.14 tokens/s and 105.89 tokens/s on a single NVIDIA A100-80GB GPU. LayerSkip (layerskip, ) proposes to exit at early layers and verify and correct with remaining layers of the model. During training, it applies layer dropout with low dropout rates for earlier layers and higher dropout rates for later layers, and adds an early exit loss to increase the accuracy of early exit at earlier layers. For the Llama2-13B model, it achieves 66.37 tokens/s on a single NVIDIA H100 GPU. LLMA (yang2023inference, ) is motivated by the observation that there are abundant identical text spans between the decoding result by an LLM and the reference that is available in many real-world scenarios (e.g., retrieved documents). LLMA first selects a text span from the reference and copies its tokens to the decoder and then efficiently checks the tokens’ appropriateness as the decoding result in parallel within one decoding step. For Llama-7B and Llama-13B models, it achieves 59.2 tokens/s and 41.1 tokens/s on a single NVIDIA V100 GPU, respectively.
@@
Skip Layer. The skip layer method (adainfer, ; RAEE, ; MOD, ) is proposed based on the idea that not all layers of LLMs are necessary during inference. AdaInfer (adainfer, ) statistically analyzes the activated layers across tasks and proposes a simple algorithm to determine the inference termination moment based on the input instance adaptively. Due to the introduction of the overhead in each layer, which is not friendly to the decoding phase, AdaInfer only takes some Benchmarks for the Q&A tasks and achieves 25.2 tokens/s for Llama2-7B on a single NVIDIA V100 GPU. Based on the basic dataflow of Adainfer, RAEE (RAEE, ) proposes to build the retrieval database to store the token information offline and leverages the information of the retrieved similar token by searching the pre-built retrieval database to guide the backbone model to exit at the layer. For the LM-BFF which is finetuned based on RoBERT-large-350M, RAEE achieves 26.23 tokens/s. The studies mentioned above both use continuous shallow layers for inference, called early exiting. MOD (MOD, ) decides whether to skip the current layer or not by pretraining the model to add a router in each layer like Mixture-of-Experts. This achieves a dynamic selection of partial layers for computation instead of forcing continuous layers. The model of MOD is not open source and no corresponding results on throughput are given in the paper.
3.3.3. ASIC
@@
Speculative Decoding. C-Transformer (kim202420, ) adopts a big-little network, which is composed of the original GPT-2 big model and a 1/10$\times$ smaller model, and a reconfigurable homogeneous architecture to increase hardware utilization and energy efficiency. During inference, only the little model computation is performed, and if the prediction probability of a specific token is over a predefined threshold, the big model computation is skipped leading to memory access reduction. Then, workloads are divided into two domains: adder-based spike domain which is efficient for small input values, and multiplier-based non-spike domain which is efficient for large input values. It has 1.431W power consumption and is fabricated in 28nm CMOS technology. For GPT-2 large model, C-Transformer achieves 2146.75 tokens/s in prefill stage.
@@
Figure 8. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with fast decoding.
3.3.4. PIM/NDP
@@
Speculative Decoding. SpecPIM (li2024specpim, ) aims to accelerate speculative inference on the PIM-enabled system by extensively exploring the heterogeneity brought by both the algorithm and the architecture. It constructs the architecture design space to satisfy each model’s disparate resource demands and dedicates the dataflow design space to fully utilize the system’s hardware resources. Based on the co-design space, it also proposes a design space exploration framework to provide the optimal design under different target scenarios. Compared with speculative inference on GPUs and existing PIM-based LLM accelerators, SpecPIM achieves 1.52$\times$/2.02$\times$ geomean speedup and 6.67$\times$/2.68$\times$ geomean higher energy efficiency.
3.3.5. Comparison
@@
In Figure 8, for fast decoding, power consumption ranges from 17.499W to 700W, with inference speeds between 24.7 tokens/s and 174.018 tokens/s. The energy efficiency ranges from 0.082 tokens/J to 9.944 tokens/J. SpecPIM (li2024specpim, ) (PIM/NDP) achieves the lowest power consumption with 17.499W, the highest throughput with batch size 1, and the highest energy efficiency.
@@
  • For GPUs, power consumption ranges from 267W to 700W, with inference speeds between 24.7 tokens/s and 169.68 tokens/s. The energy efficiency ranges from 0.082 tokens/J to 0.587 tokens/J.
@@
3.4. Operator Optimization
3.4.1. Overview
@@
Figure 9. Operator optimizations.
@@
Table 6. Operator optimization on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@
3.4.2. GPU
@@
Fusion. To address the quadratic memory requirements in the attention computation, FlashAttention (flashattention, ; flashattention2, ) fuses the attention operation into a single operator by tiling input matrices (Q, K, V) and the attention matrix into blocks. Based on FlashAttention, FlashDecoding (flashdecoding, ) proposes additionally the parallel computation along the feature dimension, improving performance for small batch size during the decoding phase. FlashDecoding achieves 95.07 tokens/s and 54.19 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. Subsequently, FlashDecoding++ (hong2024flashdecoding, ) optimizes the synchronization overhead in softmax computation by pre-determining a unified maximum based on statistical analysis in advance, enabling the parallel execution of subsequent operations and improving efficiency in typical LLMs like Llama2 and ChatGLM (chatglm, ). FlashDecoding++ achieves 115.57 tokens/s and 61.66 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. The linear operator is widely used in deep learning. In the common framework for deep learning (e.g., Pytorch), the backend of the linear operator is usually the General Matrix-Matrix Multiplication (GEMM) operation supported by NVIDIA. The naive implementation by HuggingFace achieves 44.60 tokens/s and 36.29 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. In the LLM framework (e.g., DeepSpeed (deepspeed, ), vLLM (kwon2023vllm, ), and OpenPPL (OpenPPL, )), the GEMM implementation provided by cuBLAS (cuBLAS, ) APIs is generally optimized. For the Llama2-7B model on a single NVIDIA A100 GPU, they achieve 89.24 tokens/s, 88.45 tokens/s and 93.63 tokens/s. For the Llama2-7B model on a single NVIDIA RTX 3090 GPU, they achieve 51.28 tokens/s, 53.31 tokens/s and 55.99 tokens/s. To address the inefficiency of GEMM during decoding due to the reduced dimensions, TensorRT-LLM (tensorrt-llm, ) introduces a dedicated General Matrix-Vector Multiplication (GEMV) implementation to support the decoding phase of LLM when batch size equals 1. TensorRT-LLM achieves 98.19 tokens/s for Llama2-7B on a single NVIDIA A100 GPU. For the smaller batch size during the decoding phase, FlashDecoding++ (hong2024flashdecoding, ) introduces FlatGEMM to address the inefficiencies in cuBLAS (cuBLAS, ) and CUTLASS (CUTLASS, ) libraries and employs fine-grained tiling and double buffering techniques to improve parallelism and reduce the latency of memory access. Moreover, it adopts a heuristic selection mechanism to dynamically select the most efficient operator based on the input. The performance of FlashDecoding++ is shown above. Operator fusion (flashattention, ; bytetransformer, ; deepspeed, ; hong2024flashdecoding, ) is a common and effective optimization to reduce the runtime memory access, eliminate kernel launching overhead and enhance parallelism. ByteTransformer (bytetransformer, ) and DeepSpeed (deepspeed, ) fuse the main operator (e.g., GEMM) and the lightweight operators(e.g., residual adding, layernorm and activation functions) into a single kernel to reduce the kernel launching overhead. FlashAttention (flashattention, ) mentioned above fuse the attention operator into one single kernel, reducing significantly the overhead of memory access and the memory requirements. FlashDecoding++ (hong2024flashdecoding, ) also achieves the integration of seven fused kernels in a transformer block.
3.4.3. ASIC

### v4

@@ 3.2.7 Comparison
In Figure 6, for sparsity, power consumption ranges from 33W to 450W, with inference speeds between 8.67 tokens/s and 164 tokens/s. The energy efficiency ranges from 0.06 tokens/J to 4.97 tokens/J. AccLLM [145] (FPGA) achieves the lowest power consumption with 30.0167W and the highest throughput with batch size 1 (though it also applies quantization), which also achieves the highest energy efficiency with 4.97 tokens/J.
@@
  • For GPUs, power consumption ranges from 70W to 450W, with inference speeds between 18 tokens/s and 66.69 tokens/s, situated in the middle right part of the figure. The energy efficiency ranges from 0.06 tokens/J to 0.434 token/J. Compared to CPUs, GPUs can achieve higher absolute inference speeds due to their high computing power and high bandwidth. However, the energy efficiency difference between CPUs is not significant.
  • For FPGAs, power consumption ranges from 45W to 225W, with inference speeds between 51.74 tokens/s and 164 tokens/s, also in the upper right part of the figure. The energy efficiency ranges from over 0.230 tokens/J to 4.97 tokens/J, which is much higher than GPUs and CPUs.
3.3 Fast Decoding
3.3.1 Overview
Traditional autoregressive decoding typically generates text token by token, choosing only the highest probability token at each step, known as greedy sampling. While this approach is simple and easy to implement, it may lead to a lack of diversity and creativity in the generated results. Another autoregressive method called nucleus sampling (or top-p sampling) [293] considers multiple candidates during generation by setting a cumulative probability threshold p, allowing for sampling within a certain range. Although this method offers more diversity than greedy sampling, it still operates in a step-by-step generation manner. Currently, fast decoding techniques can be mainly divided into two categories: speculative decoding and skip layer.
@@ 3.3.1 Overview
Figure 7: Fast decoding.
Table 5: Fast decoding on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@
3.3.2 CPU
ML-SpecQD [203] uses MXFP4 models as drafts. And the MXFP4 draft token generation itself can be accelerated via speculative decoding by using yet another smaller draft. This multi-Level speculative decoding with MXFP4 quantized drafts achieves speedups up to 2.72x over the BF16 baseline.
3.3.3 GPU
@@
Speculative Decoding. Speculative decoding [205] is proposed to overcome the inherently sequential process in the autoregressive decoding of LLM. The essential decoding mechanism is to make predictions (i.e., draft tokens) parallelly for multiple time steps and then select the longest prefix verified by a scoring model as the final output. Lookahead decoding [206] adopts the Guess-and-Verify paradigm as the whole decoding mechanism which generates the draft tokens by $n$-gram method and verifies the draft tokens during the forward at the same time. For Llama-7B and Llama-13B models with different datasets, it achieves 65.12-94.51 tokens/s and 56.01-90.05 tokens/s on a single NVIDIA A100-80GB GPU. To improve the acceptance rate while maintaining generation quality, many works focus on the scheme of generating draft tokens. Medusa [207] adds extra decoding heads to generate multiple subsequent tokens in parallel and uses a tree-based attention mechanism to construct multiple candidate continuations and verify them simultaneously in each decoding step. For Vicuna-7B and Vicuna-13B, it achieves 129.86 tokens/s and 98.54 tokens/s on a single NVIDIA A100-80GB GPU. EAGLE [208, 209] selects a single transformer layer with the same configuration as LLM as the draft model to make predictions autoregressively and combines the feature and token embedding as the input of the draft model. During the verification phase on the target model, EAGLE chooses the tree-based attention mechanism similar to Lookahead and Medusa to ensure the correct relationship between the draft tokens. For Vicuna-7B, Vicuna-13B, Llama2-Chat-7B, and Llama-2-Chat-13B models, it achieves 139.95 tokens/s, 132.31 tokens/s, 133.98 tokens/s, and 156.80 tokens/s on a single NVIDIA A100-80GB GPU. Based on the basic paradigm of speculative decoding for prediction by draft models and verification by target models, some studies explore the optimizations on the modules in it. Ouroboros [210] constructs a phrase candidate pool from the verification process of LLMs to provide candidates for the draft token generation of the draft model. Different from Medusa and EAGLE, Ouroboros uses the smaller LLMs (e.g., DeepSeek-7B) as the draft models for the target models (e.g., DeepSeek-34B). For Yi-34B, DeepSeek-34B and CodeLlama-34B models, it achieves 61.20 tokens/s, 41.00 tokens/s and 39.2 tokens/s on a single NVIDIA A100-80GB GPU. During the prediction phase, Sequoia [211] introduces a dynamic programming algorithm and a hardware-aware tree optimizer to find the optimal tree structure based on the runtime features and the given hardware platform. During the verification phase, Sequoia uses a novel sampling and verification method that outperforms prior work across different decoding temperatures. For Llama2-7B and Llama2-13B models, it achieves 169.68 tokens/s and 149.20 tokens/s on a single NVIDIA A100-80GB GPU.
@@
The studies mentioned above all require the help of an auxiliary model (e.g., a single transformer layer in EAGLE, several Medusa heads in Medusa, or a smaller LLM like Llama2-7B in Ouroboros) or the statistical methods (e.g., $n$-gram in Lookahead) to generate the predicted tokens and then the target model is utilized to verify the predicted tokens. Some other studies [212, 213, 214] also explore the LLM inference acceleration without the need of auxiliary models, called self-speculative decoding. Draft&Verify [212] generates draft tokens by selectively skipping certain intermediate layers of LLMs. Subsequently, the draft tokens will be verified in one forward pass. For the Llama2-13B model, it achieves 62.23 tokens/s on a single NVIDIA A100-40GB GPU. Kangaroo [213] adopts a fixed shallow sub-network of the LLM as the draft model, with the remaining layers serving as the target model. To enhance the representation ability of the draft model (i.e., the shallow sub-network), it trains an adapter module to follow the sub-network. For Vicuna-7B and Vicuna-13B models, it achieves 138.14 tokens/s and 105.89 tokens/s on a single NVIDIA A100-80GB GPU. LayerSkip [214] proposes to exit at early layers and verify and correct with remaining layers of the model. During training, it applies layer dropout with low dropout rates for earlier layers and higher dropout rates for later layers, and adds an early exit loss to increase the accuracy of early exit at earlier layers. For the Llama2-13B model, it achieves 66.37 tokens/s on a single NVIDIA H100 GPU. LLMA [204] is motivated by the observation that there are abundant identical text spans between the decoding result by an LLM and the reference that is available in many real-world scenarios (e.g., retrieved documents). LLMA first selects a text span from the reference and copies its tokens to the decoder and then efficiently checks the tokens’ appropriateness as the decoding result in parallel within one decoding step. For Llama-7B and Llama-13B models, it achieves 59.2 tokens/s and 41.1 tokens/s on a single NVIDIA V100 GPU, respectively.
AMUSD (Asynchronous Multi-device Speculative Decoding) [219] introduces a system that further accelerates generation by decoupling the draft and verify phases into a continuous, asynchronous approach. It enables both models to perform predictions independently on separate GPUs, instead of only one model(draft or verify) performing token generation at a time. For Llama3.1-8B model as a verify model and Llama3.2-1B model as a draft model, it achieves an average 71.02 tokens/s improvement over speculative decoding on two NVIDIA A100 GPUs with around 300W power consumption. PipeSpec [220] presents a framework that generalizes speculative decoding to k models arranged in a hierarchical pipeline, enabling asynchronous execution with lightweight coordination for prediction verification and rollback. For the Llama2-68M and 7B model combination, it achieves 79.18 tokens/s on a single NVIDIA A100 GPU, while for Llama3.1-1B, 8B, and 70B models, it achieves 20.51 tokens/s on two A100 GPUs. Adaptix [221] utilizes a tri-gram matrix-based LLM representation to dynamically approximate the output distribution of the LLM, allowing the model to adjust to changing token probabilities during the decoding process. It also implements a draft construction mechanism that effectively balances exploration and exploitation. For Llama2-7B and Llama-13B models, it achieves 85.62 tokens/s and 45.83 tokens/s on a NVIDIA A6000 GPU, respectively. SPIN [222] improves token speculation by using multiple heterogeneous SSMs(Small Speculative Models), with a learning-based algorithm for SSM selection that operates without prior knowledge of request difficulty. It also orchestrates speculation and verification phases by pipelining their executions on GPUs to achieve further acceleration. For Llama2-7B model as a verify model and Llama2-68M model as a draft model, it achieves 62.33 tokens/s on a NVIDIA V100 GPU. PARD [223](PARallel Draft) enables low-cost adaptation of autoregressive draft models into parallel draft models. It predicts multiple future tokens in a single forward pass of the draft phase. For Llama3.1-8B model, it achieves 274.4 tokens/s on a single NVIDIA A100 GPU. Judge Decoding [224] noticed that even GPT-4o, as a draft model, cannot achieve high acceptance rates. So it investigated how its alignment between draft and target response leads to the rejection of objectively correct continuations, and then proposed an adapted verification scheme that judges responses in a versatile way. As a result, it achieves 141.8 tokens/s with Llama3.1-8B as draft model and Llama3.1-70B as a verify model on two NVIDIA H100 GPUs. To relieve the significant latency of the draft model in speculative decoding, Falcon [225] proposes the Coupled Sequential Glancing Distillation technique to fortify inter-token dependencies within the same block during draft model training to increase speculation accuracy. For the Llama2-Chat 7B model, it achieves 147.38 tokens/s on a single NVIDIA H800 GPU.
@@
Skip Layer. The skip layer method [215, 216, 217] is proposed based on the idea that not all layers of LLMs are necessary during inference. AdaInfer [215] statistically analyzes the activated layers across tasks and proposes a simple algorithm to determine the inference termination moment based on the input instance adaptively. Due to the introduction of the overhead in each layer, which is not friendly to the decoding phase, AdaInfer only takes some Benchmarks for the Q&A tasks and achieves 25.2 tokens/s for Llama2-7B on a single NVIDIA V100 GPU. Based on the basic dataflow of Adainfer, RAEE [216] proposes to build the retrieval database to store the token information offline and leverages the information of the retrieved similar token by searching the pre-built retrieval database to guide the backbone model to exit at the layer. For the LM-BFF which is finetuned based on RoBERT-large-350M, RAEE achieves 26.23 tokens/s. The studies mentioned above both use continuous shallow layers for inference, called early exiting. MOD [217] decides whether to skip the current layer or not by pretraining the model to add a router in each layer like Mixture-of-Experts. This achieves a dynamic selection of partial layers for computation instead of forcing continuous layers. The model of MOD is not open source and no corresponding results on throughput are given in the paper. SpecEE [218] predicts whether it is early existing through a lightweight predictor by reducing the vocabulary space. And for this method, it designs the Two-level heuristic predictor scheduling and the Context-aware merged mapping for the predictor. For Llama-7B and Llama-13B models, SpecEE + EAGLE achieves 124.66 tokens/s and 120.8 tokens/s on a single NVIDIA A100 GPU, respectively.
3.3.4 ASIC
@@
Speculative Decoding. C-Transformer [226] adopts a big-little network, which is composed of the original GPT-2 big model and a 1/10$\times$ smaller model, and a reconfigurable homogeneous architecture to increase hardware utilization and energy efficiency. During inference, only the little model computation is performed, and if the prediction probability of a specific token is over a predefined threshold, the big model computation is skipped leading to memory access reduction. Then, workloads are divided into two domains: adder-based spike domain which is efficient for small input values, and multiplier-based non-spike domain which is efficient for large input values. It has 1.431W power consumption and is fabricated in 28nm CMOS technology. For GPT-2 large model, C-Transformer achieves 2146.75 tokens/s in prefill stage.
@@
Figure 8: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with fast decoding.
3.3.5 PIM/NDP
@@
Speculative Decoding. SpecPIM [227] aims to accelerate speculative inference on the PIM-enabled system by extensively exploring the heterogeneity brought by both the algorithm and the architecture. It constructs the architecture design space to satisfy each model’s disparate resource demands and dedicates the dataflow design space to fully utilize the system’s hardware resources. Based on the co-design space, it also proposes a design space exploration framework to provide the optimal design under different target scenarios. Compared with speculative inference on GPUs and existing PIM-based LLM accelerators, SpecPIM achieves 1.52$\times$/2.02$\times$ geomean speedup and 6.67$\times$/2.68$\times$ geomean higher energy efficiency.
3.3.6 Comparison
@@
In Figure 8, for fast decoding, power consumption ranges from 17.499W to 700W, with inference speeds between 13.7 tokens/s and 274.4 tokens/s. The energy efficiency ranges from 0.082 tokens/J to 9.944 tokens/J. SpecPIM [227] (PIM/NDP) achieves the lowest power consumption with 17.499W, the highest throughput with batch size 1, and the highest energy efficiency.
@@
  • For GPUs, power consumption ranges from 85W to 700W, with inference speeds between 13.7 tokens/s and 274.4 tokens/s. The energy efficiency ranges from 0.082 tokens/J to 0.686 tokens/J.
@@
3.4 Operator Optimization
3.4.1 Overview
@@
Figure 9: Operator optimizations.
@@
Table 6: Operator optimization on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@
3.4.2 CPU
Storage Optimization. For the on-device inference of LLM with resource constraints, FlexInfer [229] proposes methods like asynchronous prefetching, balanced memory locking, and flexible tensor preservation to enhance memory efficiency and relieve bandwidth bottlenecks. For the Llama2-7B model, it achieves 12 tokens/s on an AMD 7995WX CPU.
Fusion. V-Seek [228] optimizes LLM inference on the Sophon SG2042, the first commercially available many-core RISC-V CPU with vector processing capabilities. For DeepSeek R1 Distill Llama 8B and DeepSeek R1 Distill QWEN 14B, we achieve 4.32/2.29 token/s for token generation and 6.54/3.68 token/s for prompt processing.
3.4.3 GPU
@@
Fusion. To address the quadratic memory requirements in the attention computation, FlashAttention [230, 231] fuses the attention operation into a single operator by tiling input matrices (Q, K, V) and the attention matrix into blocks. Based on FlashAttention, FlashDecoding [232] proposes additionally the parallel computation along the feature dimension, improving performance for small batch size during the decoding phase. FlashDecoding achieves 95.07 tokens/s and 54.19 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. Subsequently, FlashDecoding++ [233] optimizes the synchronization overhead in softmax computation by pre-determining a unified maximum based on statistical analysis in advance, enabling the parallel execution of subsequent operations and improving efficiency in typical LLMs like Llama2 and ChatGLM [33]. FlashDecoding++ achieves 115.57 tokens/s and 61.66 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. The linear operator is widely used in deep learning. In the common framework for deep learning (e.g., Pytorch), the backend of the linear operator is usually the General Matrix-Matrix Multiplication (GEMM) operation supported by NVIDIA. The naive implementation by HuggingFace achieves 44.60 tokens/s and 36.29 tokens/s for Llama2-7B on a single NVIDIA A100 and RTX 3090 GPU respectively. In the LLM framework (e.g., DeepSpeed [234], vLLM [235], and OpenPPL [236]), the GEMM implementation provided by cuBLAS [237] APIs is generally optimized. For the Llama2-7B model on a single NVIDIA A100 GPU, they achieve 89.24 tokens/s, 88.45 tokens/s and 93.63 tokens/s. For the Llama2-7B model on a single NVIDIA RTX 3090 GPU, they achieve 51.28 tokens/s, 53.31 tokens/s and 55.99 tokens/s. To address the inefficiency of GEMM during decoding due to the reduced dimensions, TensorRT-LLM [238] introduces a dedicated General Matrix-Vector Multiplication (GEMV) implementation to support the decoding phase of LLM when batch size equals 1. TensorRT-LLM achieves 98.19 tokens/s for Llama2-7B on a single NVIDIA A100 GPU. For the smaller batch size during the decoding phase, FlashDecoding++ [233] introduces FlatGEMM to address the inefficiencies in cuBLAS [237] and CUTLASS [239] libraries and employs fine-grained tiling and double buffering techniques to improve parallelism and reduce the latency of memory access. Moreover, it adopts a heuristic selection mechanism to dynamically select the most efficient operator based on the input. The performance of FlashDecoding++ is shown above. Operator fusion [230, 240, 234, 233] is a common and effective optimization to reduce the runtime memory access, eliminate kernel launching overhead and enhance parallelism. ByteTransformer [240] and DeepSpeed [234] fuse the main operator (e.g., GEMM) and the lightweight operators(e.g., residual adding, layernorm and activation functions) into a single kernel to reduce the kernel launching overhead. FlashAttention [230] mentioned above fuse the attention operator into one single kernel, reducing significantly the overhead of memory access and the memory requirements. FlashDecoding++ [233] also achieves the integration of seven fused kernels in a transformer block.
SpInfer [191] is a high-performance framework tailored for sparsified LLM inference on GPUs. It introduces Tensor-Core-Aware Bitmap Encoding (TCA-BME) to minimize indexing overhead and integrates an optimized SpMM kernel with Shared Memory Bitmap Decoding (SMBD) and an asynchronous pipeline. It achieves around 500 tokens/s with batch size 8 for OPT-13B model on a single RTX 4090 GPU.
Storage Optimization. FlashFormer [241] introduces a proof-of-concept kernel for accelerating single-batch inference, which exploits increased overlapping of memory movement and amortizes the launch overhead for the kernel over the entire forward pass. For Llama3.1-8B model, it achieves 181 tokens/s on a NVIDIA H100 GPU. To tackle the memory bandwidth bottleneck during LLM inference, Dong et al. [242] proposes L2 Cache-oriented asynchronous KV Cache prefetching method through computation-load overlap, which achieves around 3000 tokens/s for the Llama2-7B model on a single NVIDIA H20 GPU with batch size 64.
3.4.4 FPGA
Nonlinear Function Approximation. HAAN [243] is a holistic normalization acceleration approach for LLMs. HAAN exploits the correlation in normalization statistics among adjacent layers to bypass normalization computation by estimating statistics from preceding layers.
3.4.5 ASIC